OpenStudy (janu16):

Three functions are given below: f(x), g(x), and h(x). Explain how to find the axis of symmetry for each function, and rank the functions based on their axis of symmetry (from smallest to largest). f(x) = –4(x − 8)2 + 3 g(x) = 3x2 + 12x + 15

OpenStudy (janu16):

This

OpenStudy (janu16):

@pooja195

OpenStudy (janu16):

@Nnesha

Nnesha (nnesha):

\[\huge\rm y=a(x-\color{red}{h})^2+\color{blue}{k}\] vertex form of quadratic equation where (h,k) is the vertex and x-coordinate of the vertex represent x-axis of symmetry axis of symm. = h

OpenStudy (janu16):

ok.. so how do you find it like plug in this formula?

Nnesha (nnesha):

if the equation is already in vertex form then you don't need to plugin anything just look at the equation what number is replaced by h ? if the given equation is standard then u can use this formula to find x-coordinate of the vertex \[\large\rm x=-\frac{b}{2a}\] b= coefficient of x term a=leading coefficient \[\rm y=ax^2+bx+c\]

Nnesha (nnesha):

standard form*

OpenStudy (janu16):

f(x) = –4(x − 8)2 + 3 8 is replaced by h

OpenStudy (janu16):

so axis of symmetry for f(x) would be 8?

Nnesha (nnesha):

yes right.

OpenStudy (janu16):

for g(x) do you have to plug in the formula?

Nnesha (nnesha):

yes right

OpenStudy (janu16):

you just rreplace the numbers or you have to solve it or something?

Nnesha (nnesha):

first replace the number with their values and then simplify

OpenStudy (janu16):

would it be g(x)= 3(x-12)^2 +15?

Nnesha (nnesha):

hmm how did convert g(x) standard form to vertex ? are you familiar with the `completing the square ` method ?

OpenStudy (janu16):

yaa

Nnesha (nnesha):

the formula i gave you is simple and easier than completing the square it will take just few minutes but it's okay if you want to complete the square

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Janu16 would it be g(x)= 3(x-12)^2 +15? \(\color{blue}{\text{End of Quote}}\) that's not correct show me the work how did you get that equation :=))

OpenStudy (janu16):

ohh i just replaced the numbers. I didn't simplify

Nnesha (nnesha):

no you can't just replace number to arrive from \[\rm Ax^2+Bx+C \] to \[\rm a(x-h)^2+k\] you have to apply completing the square method you can just replace the numbers around

Nnesha (nnesha):

use this formula to find x-coordinate of vertex (which is equal to axis of symm)\[\rm x=-\frac{b}{2a}\] replace b and a in this formula by their values

OpenStudy (janu16):

i don't know how to use that formula. if you don't mind can you help solve this with completing the sqaure cause thats what im learning

Nnesha (nnesha):

ohh sure! for that formula we just have to replace b with 12 and a with 3 \[\large\rm y=\color{reD}{a}x^2+\color{blue}{b}x+\color{green}{c}\]\[\large\rm y=\color{reD}{3}x^2+\color{blue}{12}x+\color{green}{15}\] a=leading coefficient b=middle term( coefficient of x term) c=constant term

Nnesha (nnesha):

let's complete the square \[\large\rm g(x)=3x^2+12x+15\] since the leading coefficient isn't one it would be great if we first take out the GCF(greatest common factor ) of first two terms

Nnesha (nnesha):

\[\rm g(x) = \color{Red}{(3x^2+12x)}+15\] we aren't gonna touch the constant term till the end

Nnesha (nnesha):

what's the greatest common factor in first two terms ?

OpenStudy (janu16):

4

OpenStudy (janu16):

wait no

OpenStudy (janu16):

3

Nnesha (nnesha):

right take out 3 from (3x^2 +12x) what would u have left with ?

OpenStudy (janu16):

x+4

Nnesha (nnesha):

hmm just take out the common factor 3 not the x variable the reason is the to get the 1 for leading coefficient so it would be \[\large\rm g(x) =3(\color{reD}{x^2+4x})+15\]now we are going to complete the square of `x^2+4x` 1st) we should take `half of x term` and then square it

OpenStudy (janu16):

half of 4 would be 2

Nnesha (nnesha):

\[\huge\rm y=(\color{ReD}{x^2+bx})+c\] i would take of of x term which is b in this example and then subtract of (b/2)^2 from constant term \[\large\rm y=(\color{reD}{x^2+bx} +(\frac{b}{2})^2)+c - (\frac{b}{2})^2\]

Nnesha (nnesha):

remember the factors of this quadratic equation would always be \[\rm (x+\frac{b}{2})^2\] (x+ half of b term)^2

Nnesha (nnesha):

\[\large\rm y=(\color{blue}{x^2+bx +(\frac{b}{2})^2})+c - (\frac{b}{2})^2\] \[\large\rm y=\color{blue}{(x+\frac{b}{2})^2} +c -(\frac{b}{2})^2\]

OpenStudy (janu16):

so\[(x^2+4x)15-4\]

OpenStudy (janu16):

is that correct^^

OpenStudy (janu16):

or is it (x2+4x + 4)15−4

Nnesha (nnesha):

well there is something missing you should add (b/2)^2 in the parentheses to make it trinomial equation

Nnesha (nnesha):

that looks good but don't forget the sign +15-4

Nnesha (nnesha):

hmm but wait we have common factor at front og the parentheses

Nnesha (nnesha):

\[\large\rm g(x) =3(\color{reD}{x^2+4x+4})+15\] when we have a GCF we should multiply (b/2)^2 by gcf and THEN subtract from constant term

OpenStudy (janu16):

so 4 times 4?

Nnesha (nnesha):

no (b/2)^2 = ??

OpenStudy (janu16):

4

OpenStudy (janu16):

8

Nnesha (nnesha):

which one ?? :=)) b is what ?

OpenStudy (janu16):

so you do 4 divided by 2 which is 2 than sqaure it which is 4

Nnesha (nnesha):

yes right 4 is correct 8 isn't now multiply 4 times 3(gcf) then subtract from the constant term

OpenStudy (janu16):

so you 4 times 3 which is 12

OpenStudy (janu16):

than 15-12?

Nnesha (nnesha):

right

Nnesha (nnesha):

so what would be the final equation ?

OpenStudy (janu16):

\[x^2+4x=3\]

Nnesha (nnesha):

hmm no

OpenStudy (janu16):

we dont write the 3 in the beginning because we got rid of it right?

Nnesha (nnesha):

we didn't get rid of it just multiply it by (b/2)^2 to subtract it from constant term

OpenStudy (janu16):

x^2+4x+3=0

Nnesha (nnesha):

not really we are working on one side so \[\large\rm g(x) =3(\color{reD}{x^2+4x+4})+15\color{blue}{-12}\] remember we got the trinomial equation

OpenStudy (janu16):

ok..

Nnesha (nnesha):

and the factor of that trinomial would always be \[\rm (x+\frac{b}{2})^2\]

OpenStudy (janu16):

so what would the equation look like if you multiply 4 and 3 and subtract 15 from 12?

Nnesha (nnesha):

15-12= 3 right now deal with this equation (x^2+4x+4) what are the factors ?

Nnesha (nnesha):

like i said when we complete teh square the factors of the trinomial would always be \[\rm (x+\frac{b}{2})^2\] b= coefficient of x term

Nnesha (nnesha):

does it make sense ?

OpenStudy (janu16):

yaa

Nnesha (nnesha):

alright then what would be the factor of (x^2+4x+4) ?

OpenStudy (janu16):

lie factoring the whole thing?

OpenStudy (janu16):

x+2^2

Nnesha (nnesha):

factor the trinomial (quadratic equation )(red part in this equatin) \[\large\rm g(x) =3(\color{reD}{x^2+4x+4})+15\color{blue}{-12}\] ( quadratic equation )

Nnesha (nnesha):

not just 2 to the 2 power

Nnesha (nnesha):

remember u will always get \[\rm (x+\frac{b}{2})^2\] in that equation b is 4 so \[(x+\frac{4}{2})^2 \rightarrow (x+2)^2\]

Nnesha (nnesha):

so what would be the final equation ?

Nnesha (nnesha):

for h(x) remember